What is a point that goes through the line y=2/3x+3
Found 2 solutions by stanbon, funmath:
Answer by stanbon(75887)

What is a point that goes through the line y=2/3x+3
Well, points don't go through lines but (0,3)
is on the line.
Cheers,
Stan H.

Answer by funmath(2933)

There are an infinite number of points on a line. The easiest one to find in this case is the y-intercept. Let x=0 and solve for y.
What is a point that goes through the line y=2/3x+3
y=(2/3)(0)+3
y=0+3
y=3
The point (x,y)=(0,3) is on the line.
You can let x be anything you want, but if you want integers for y, let x be things divisible by 3.
When x=-3
y=(2/3)(-3)+3
y=-6/3+3
y=-2+3
y=1
The point (x,y)=(-3,1) is also on the line.
Why don't you find the point when x=3, and tell me what you get?
Happy Calculating!!!